A THEORETICAL RE-EXAMINATION OF THE UV PHOTODISSOCIATION SPECTRA OF HI AND DI

\underline{ROBERT J. LE ROY} and GEOFFREY T. KRAEMER, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

Since the detailed theoretical study Levy and Shapiro^{a} over a decade ago, incisive new experimental studies have been reported which contradict older measurements of the anisotropy parameters associated with the I*(^{2}P_{1/2}) dissociation channels, but disagree among themselves regarding the magnitude of the branching ratio between the I(^{2}P_{3/2}) and I*(^{2}P_{1/2}) channels.^{b}^{,}^{c} Previous theoretical studies have also tended to focus on HI, and inconsistently model or overlook apparent discrepancies with the total photodissociation cross-sections for DI. This paper reports our progress on a theoretical re-examination of this system which uses direct least-squares fits to all of the various types of data to attempt to delineate better the underlying potential energy curves and transition moment functions.

a I. Levy and M. Shapiro, J. Chem. Phys. 89, 2900 (1988)

b S.R. Langford, P.M. Regan, A.J. Orr-Ewing and M.N.R. Ashfold, Chem. Phys. 231, 245 (1998)

c D. Gendron and J.W. Hepburn, J. Chem. Phys. 109, 7204 (1998)